OpenStudy (anonymous):

Simplify: (4x − 6) + (3x + 6)

OpenStudy (hoblos):

(4x − 6) + (3x + 6) = 4x − 6 + 3x + 6 = 7x

OpenStudy (anonymous):

that is a slight simplifification but ill just do it my self im just being lazy o my home work

OpenStudy (anonymous):

=4x-6+3x+6 =4x+3x+6-6 =7x+0 =7x